OpenStudy (anonymous):

A bicycle tire has a diameter of 20 inches and is revolving at a rate of 10 rpm. At t = 0, a certain point is at height 0. What is the height of the point above the ground after 20 seconds? I need the formula for this and I'll figure out the rest c:

OpenStudy (tkhunny):

The formula is the important part. 20 seconds = 1/3 minute 10 rpm for 1/3 minute is 3 + 1/3 revolutions

OpenStudy (agent0smith):

 height above ground = 20+ h inches

OpenStudy (anonymous):

Oh okai thanks guys c:

OpenStudy (agent0smith):

You can find angle x with tkhunny's info, what's 1/3rd of a circle in degrees? That amount will equal x+90 degrees. Then use sin x = h/20.

OpenStudy (anonymous):

the h equals 1 correct?

OpenStudy (anonymous):

I mean does the h = the sin x?

OpenStudy (agent0smith):

sin x = h/20 so 20*sin x = h

OpenStudy (anonymous):

ohhh okai c:

OpenStudy (agent0smith):

Did you find angle x as well?

OpenStudy (agent0smith):

Don't forget, once you find h, the height above ground = 20+ h inches

OpenStudy (anonymous):

oh okai I didn't do that last part and yeah I found out my x value